# Covariance test and universal bootstrap by operator norm

### v1

@@ 2 Proposed method and theoretical outlines
The empirical distribution of $T^{\text{ub},b}$ for $b=1,\dots,B$ serves as an approximation for the distribution of $T$. In particular, we use the empirical upper-$\alpha$ quantile, $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}_{0}}(\alpha)$, of $T^{\text{ub},b}$, $b=1,\dots,B$, as the threshold for the test (5) based on $T$. Defining $\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)=\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0},\bm{\Sigma}_{0}}(\alpha)$, we reject $H_{0}$ if $T\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)$. Results like (7) ensure the consistency of $\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)$ to the upper-$\alpha$ quantile $q_{\bm{\Sigma}_{0}}(\alpha)$ of $T$ under $H_{0}$ as $n\to\infty$ and $B\to\infty$, validating the universal bootstrap procedure.
where $T$ is a statistic, $T^{*}$ is its Gaussian counterpart, and $\mathcal{A}$ represents a specified family of sets. For example, Chernozhukov, Chetverikov and Kato (2013, 2017); Chernozhukov, Chetverikov and Koike (2023) considered $T$ as mean estimators and $\mathcal{A}$ as all rectangular in $\mathbb{R}^{p}$. A more related choice of $\mathcal{A}$ is in Zhai (2018); Xu, Zhang and Wu (2019); Fang and Koike (2024), who also considered mean estimators but take $\mathcal{A}$ to be the sets of Euclidean balls and convex sets in $\mathbb{R}^{p}$. Their results demonstrated that under mild conditions and for sets of Euclidean balls $\mathcal{A}$, $\rho(\mathcal{A})$ converge to 00 if and only if $p/n\to 0$, meaning the Gaussian approximation holds when $p=o(n)$. For comparison, we observe that the operator norm ball $\mathbf{B}_{\text{op}}(\bm{v},t)$ for vector $\bm{v}$ in $\mathbb{R}^{p}$ coincides with the Euclidean balls, and our results show that the universality approximation holds when $p/n$ converges to a nonzero constant or even diverges to infinity. For the covariance test, Han, Xu and Zhou (2018) took $T$ as the sample covariance matrix and $\mathcal{A}$ as all sets of $s$-sparse operator norm balls (defined in their work). Especially, with $\mathcal{A}$ as all operator norm balls, i.e. $s=p$, they required $p=o(n^{1/9})$, limiting $p$ to be considerably smaller than $n$. Similarly, Lopes (2022a); Lopes, Erichson and Mahoney (2023) considered $T$ as sample covariance with $\mathcal{A}$ as operator norm balls, but imposed a decay rate $i^{-\beta}$ for the $i$-th largest eigenvalue $\lambda_{i}(\bm{\Sigma})$ of $\bm{\Sigma}$ with $\beta>1$, implying a low intrinsic test dimension. Likewise, Giessing (2023) required the effective rank $r(\bm{\Sigma})=\text{tr}(\bm{\Sigma})/|\bm{\Sigma}|_{\text{op}}$ to satisfy $r(\bm{\Sigma})=o(n^{1/6})$. In contrast, we impose no such assumptions, allowing each eigenvalue of $\bm{\Sigma}$ to be of comparable scale. To summarize, previous work has typically assumed either $p\ll n$ or fast-decaying eigenvalues, yielding consistent estimates for $\bm{\Sigma}$. In our setting, however, no consistent estimator of $\bm{\Sigma}$ exists (see Cai and Ma (2013)). These comparisons underscore the advantages of our proposal, even in regimes lacking consistency.
Given these improvements on existing results, we establish a universal result that extends beyond (7).
Result 2 (Informal).
For $\bm{\Sigma}_{0}$ that commutes with $\bm{\Sigma}$, we have
$\rho_{n}(\bm{\Sigma}_{0})\leq Cn^{-\delta}\to 0\quad\text{as}\quad n\to\infty.$ (8)
for some constants $C$ and $\delta>0$. See Theorem 3.5 for a formal description.
This result generalizes the universal bootstrap consistency of $T$ under $H_{0}$ to alternative covariance $\bm{\Sigma}_{0}$ distinct from $\bm{\Sigma}$. The commutativity requirement between $\bm{\Sigma}_{0}$ and $\bm{\Sigma}$ assumes the two matrices share eigenvectors. Similar assumptions appear in Zhou, Bai and Hu (2023); Zhou et al. (2024). The result in (8) further guarantees universal bootstrap consistency for statistics beyond $T$. For instance, we can estimate $\bm{\Sigma}$ with shrinkage estimators such as $a\hat{\bm{\Sigma}}+(1-a)\bm{I}_{p}$ for $a\in(0,1)$ as proposed by Schäfer and Strimmer (2005), or $a\hat{\bm{\Sigma}}$ for some $a>0$ as in Tsukuma (2016). Using these estimators, covariance can be tested with statistics $T^{\text{shr}}_{1}=\left|a\hat{\bm{\Sigma}}+(1-a)\bm{I}_{p}-\bm{\Sigma}_{0}\right|_{\text{op}}=a\left|\hat{\bm{\Sigma}}-(\bm{\Sigma}_{0}-(1-a)\bm{I}_{p})/a\right|_{\text{op}}$ or $T^{\text{shr}}_{2}=\left|a\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}\right|_{\text{op}}=a\left|\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}/a\right|_{\text{op}}$. Under $H_{0}$, universal bootstrap consistency for $T^{\text{shr}}_{1}$ and $T^{\text{shr}}_{2}$ can be shown by $\rho_{n}\left((\bm{\Sigma}_{0}-(1-a)\bm{I}_{p})/a\right)\to 0$ and $\rho_{n}\left(\bm{\Sigma}_{0}/a\right)\to 0$, as $(\bm{\Sigma}_{0}-(1-a)\bm{I}_{p})/a$ and $\bm{\Sigma}_{0}/a$ commute with $\bm{\Sigma}$.
Despite the broad applications of (8), certain statistics using extreme eigenvalues remain outside its scope. For example, as outlined in Section 1, we define $T^{\text{Com}}$ to combine $T$ and $T^{\text{Roy}}$ for enhanced statistical power, where $T^{\text{Com}}$ is given by
@@
While (8) demonstrates universal bootstrap consistency for $T$ and $T^{\text{Roy}}$ separately, it does not apply to $T^{\text{Com}}$. This limitation arises from $T^{\text{Com}}$’s dependence on the joint law of $\sigma_{1}\left(\bm{\hat{\Sigma}}-\bm{\Sigma}_{0}\right)$ and $\sigma_{1}\left(\bm{\Sigma}_{0}^{-1/2}\bm{\hat{\Sigma}}\bm{\Sigma}_{0}^{-1/2}-\bm{I}_{p}\right)$, which introduces a complex dependence structure. To address these limitations, we develop a generalized universality theorem that accounts for dependencies among various extreme eigenvalues. We begin by introducing the generalized operator norm ball $\mathbf{B}_{k,\text{op}}(\bm{\Sigma}_{1},\bm{\Sigma}_{2},\bm{t}_{k})$
where $\bm{\Sigma}_{2}$ is positive-definite, and $\bm{t}_{k}=(t_{1},\cdots,t_{k})^{T}$ with $t_{i}\geq 0$ for each $i=1,\cdots,k$. Since the operator norm $\|\bm{N}-\bm{\Sigma}_{1}\|_{\text{op}}$ equals the largest singular value $\sigma_{1}(\bm{N}-\bm{\Sigma}_{1})$, we obtain $\mathbf{B}_{\text{op}}(\bm{\Sigma},t)=\mathbf{B}_{1,\text{op}}(\bm{\Sigma},\bm{I}_{p},t)$. Next, we define the following quantity for symmetric matrices $\bm{\Sigma}_{1,m}$ positive-definite matrices $\bm{\Sigma}_{2,m}$, $m=1,\cdots,M$,
Result 3 (Informal).
For $\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}_{1,m}\bm{\Sigma}_{2,m}^{-1/2}$ that commutes with $\bm{\Sigma}_{2,m}^{-1/2}\bm{\Sigma}\bm{\Sigma}_{2,m}^{-1/2}$ for $m=1,\cdots,M$, we have
$\rho_{n}(\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M})\leq Cn^{-\delta}\to 0\quad\text{as}\quad n\to\infty.$ (12)
Take $M=1$, $k_{1}=1$, $\bm{\Sigma}_{1,1}=\bm{\Sigma}_{0}$, $\bm{\Sigma}_{2,1}=\bm{I}_{p}$, (12) recovers (8). But (12) provides a more general result, showing that the joint law of
@@
$\left\{\left(\sigma_{k}\left(\bm{\Sigma}_{2,1}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,1})\bm{\Sigma}_{2,1}^{-\frac{1}{2}}\right)\right)_{k=1}^{k_{1}},\cdots,\left(\sigma_{k}\left(\bm{\Sigma}_{2,M}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,M})\bm{\Sigma}_{2,M}^{-\frac{1}{2}}\right)\right)_{k=1}^{k_{M}}\right\},$ (13)
can be approximated by the universal bootstrap. This result allows for constructing statistics by combining extreme eigenvalues in (13) in various ways, with (12) confirming universal bootstrap consistency for these statistics.
3 Universal bootstrap
3.1 Preliminaries
Our results rely on the universality from the random matrix theory. To proceed, we first introduce some preliminary results relevant to our analysis. Denote $\bm{X}_{i}=\bm{\Sigma}^{1/2}\bm{Z}_{i}$, where $\bm{Z}_{i}$ has zero mean and identity covariance matrix for $i=1,\cdots,n$. We accordingly define $\bm{Z}=(\bm{Z}_{1},\cdots,\bm{Z}_{n})^{T}\in\mathbb{R}^{n\times p}$. The primary matrix of interest is $\bm{\hat{\Sigma}}-\bm{\Sigma}_{0}$. This inspires us to consider the matrix of a general form
$\bm{M}_{n}=\bm{\hat{\Sigma}}+\bm{R}=\phi^{\frac{1}{2}}\bm{M}^{\prime}_{n}$ (14)
where $\bm{M}^{\prime}_{n}=(np)^{-1/2}\ \bm{\Sigma}^{1/2}\bm{Z}^{T}\bm{Z}\bm{\Sigma}^{1/2}+\bm{R}^{\prime}$ and $\bm{R}^{\prime}=\phi^{-1/2}\ \bm{R}$. Here, $\bm{R}$ is a symmetric matrix with eigenvalues that may be positive, negative, or zero, and $\bm{M}^{\prime}$ is normalized to address cases where $\alpha>1$, as per Ding and Wang (2023). We impose some assumptions to determine the limit of the Stieltjes transform $m_{\bm{M}^{\prime}_{n}}(z)=\text{tr}(\bm{M}_{n}^{\prime}-z\bm{I}_{p})^{-1}/p$ of $\bm{M}^{\prime}_{n}$.
@@ 3.1 Preliminaries
Suppose that $\bm{Z}=(Z_{ij})_{n\times p}$ and $\left\{Z_{ij},i=1,\cdots,n;j=1,\cdots,p\right\}$ are independent with $\mathbb{E}[Z_{ij}]=0$, $\mathbb{E}[Z_{ij}^{2}]=1$. There exists a positive sequence $C_{k}$ such that $\mathbb{E}[|Z_{ij}|^{k}]\leq C_{k}$ for $i=1,\cdots,n;j=1,\cdots,p$ and $k\in\mathbb{N}$.
@@
Assumptions 1 and 2 are standard and often appear in the random matrix literature. Notably, Assumption 1 relies only on independence, without assuming identical distribution as in Qiu, Li and Yao (2023). While all moments exist under Assumption 1, this condition could be relaxed as discussed in Ding and Yang (2018). We do not pursue this here. We also permit $\bm{\Sigma}$ to be singular, a less restrictive condition than the invertibility assumed in Ding and Wang (2023); Ding, Hu and Wang (2024). This generalization enables testing singular $\bm{\Sigma}_{0}$, where $T^{\text{Roy}}$ is invalid but $T$ remains applicable. Lastly, Assumption 3, imposed for technical requirements, necessitates that $\bm{\Sigma}$ and $\bm{R}$ share the same eigenvectors. The same Assumption is required in the signal-plus-noise model as in Zhou, Bai and Hu (2023); Zhou et al. (2024).
We introduce the following deterministic equivalence matrix
$\bar{\bm{Q}}_{p}(z)=\bigg{(}\bm{R}^{\prime}+\frac{1}{\phi^{\frac{1}{2}}(1+\phi^{\frac{1}{2}}e(z))}\bm{\Sigma}-z\bm{I}_{p}\bigg{)}^{-1},$ (15)
where $\phi=p/n$ and $e(z)$ is the fixed point of the equation $e(z)=\text{tr}\big{(}\bm{R}^{\prime}+\phi^{-1/2}(1+\phi^{1/2}e(z))^{-1}\bm{\Sigma}-z\bm{I}_{p}\big{)}^{-1}\bm{\Sigma}/p$. Using $\bar{\bm{Q}}_{p}(z)$, we define the associated Stieltjes transform $\bar{m}_{p}(z)$ as $\bar{m}_{p}(z)=\text{tr}(\bar{\bm{Q}}_{p}(z))/p$. According to Couillet, Debbah and Silverstein (2011), $e(z)$ and $\bar{\bm{Q}}_{p}(z)$ are well-defined for every $z\in\mathbb{C}^{+}$, and $\bar{m}_{p}(z)$ serves as the Stieltjes transform of a measure $\rho$ on $\mathbb{R}$. We further denote $E_{+}$ and $E_{-}$ as the endpoints of $\rho$. Formally, if we define $\text{supp}(\rho)=\left\{x\in\mathbb{R}\ :\ \rho([x-\epsilon,x+\epsilon])>0,\ \forall\epsilon>0\right\}$, we have $E_{+}=\sup\text{supp}(\rho)$ and $E_{-}=\inf\text{supp}(\rho)$. Combining results of Knowles and Yin (2017) for the $\alpha=1$ case and Ding and Wang (2023) for the $\alpha>1$ case, it follows that $E_{+}\asymp\phi^{1/2}$ and $|E_{-}|\lesssim 1$. Intuitively, the largest eigenvalue $\lambda_{1}(\bm{M}^{\prime}_{n})$ approaches $E_{+}$, while $\lambda_{p}(\bm{M}^{\prime}_{n})$ approaches $E_{-}$. The next subsection characterizes the fluctuations of $\lambda_{1}(\bm{M}^{\prime}_{n})-E_{+}$ and $\lambda_{p}(\bm{M}^{\prime}_{n})-E_{-}$.
3.2 Universality
In this subsection, we establish the anisotropic local law, which forms the foundation for our universality results. Specifically, we aim to describe the fluctuations of $\lambda_{1}(\bm{M}^{\prime}_{n})-E_{+}$ and $\lambda_{p}(\bm{M}^{\prime}_{n})-E_{-}$. This requires us to analyze the convergence of the Stieltjes transform near the endpoints $E_{+}$ and $E_{-}$. We define the local domains $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ as
$\mathcal{D}_{\pm}=\mathcal{D}_{\pm}(\tau)=\left\{z=E+\text{i}\eta\in\mathbb{C}^{+}\ :\ |z|\geq\tau,|E-E_{\pm}|\leq\tau^{-1},n^{-1+\tau}\leq\eta\leq\tau^{-1}\right\}$ (16)
for a fixed parameter $10<\tau<$ 1. For $\alpha=1$, let $\mathcal{D}=\mathcal{D}_{+}\cup\mathcal{D}_{-}$, and for $\alpha>1$, define $\mathcal{D}=\mathcal{D}_{+}$. We will focus on the behavior of the Green function on $\mathcal{D}$. This definition ensures that for $\alpha=1$, both the largest and smallest eigenvalues of $\bm{M}_{n}^{\prime}$ are controlled, so that $\sigma_{1}(\bm{M}_{n}^{\prime})=\max\{|\lambda_{1}(\bm{M}_{n}^{\prime})|,|\lambda_{p}(\bm{M}_{n}^{\prime})|\}$ can also be controlled. While for the case $\alpha>1$, we always have $|E_{+}|\gg|E_{-}|$, thus only the largest eigenvalue requires attention.
To facilitate the presentation of the anisotropic local law, we assume in this subsection that
[…]
Under (17), we can rewrite $\bm{M}^{\prime}_{n}$ as $\bm{M}^{\prime}_{n}=\bm{\Sigma}^{1/2}\left((np)^{-1/2}\ \bm{Z}^{T}\bm{Z}+\bm{R}^{{}^{\prime\prime}}\right)\bm{\Sigma}^{1/2}$ where $\bm{R}^{{}^{\prime\prime}}=\bm{\Sigma}^{-1/2}\bm{R}^{{}^{\prime}}\bm{\Sigma}^{-1/2}$. Since $\bm{M}^{\prime}_{n}$ is quadratic in $\bm{X}$, we follow Knowles and Yin (2017) to define the linearization matrix $\bm{H}(z)$ and the corresponding linearized Green function $\bm{G}(z)$
$\bm{H}(z)=\left[\begin{array}[]{ccc}-\bm{\Sigma}^{-1}&(\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}&\frac{1}{(np)^{\frac{1}{4}}}\bm{Z}^{T}\\ (\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}&-z\bm{I}_{p}&\bm{0}\\ \frac{1}{(np)^{\frac{1}{4}}}\bm{Z}&\bm{0}&-z\bm{I}_{n}\end{array}\right],\quad\bm{G}(z)=\bm{H}^{-1}(z).$ (21)
Notice that $\bm{G}(z)$ is a large $(n+2p)\times(n+2p)$ matrix. We also define the deterministic equivalent Green function as
$\bar{\bm{G}}(z)=\left[\begin{array}[]{ccc}z\bm{\Sigma}^{\frac{1}{2}}\bar{\bm{Q}}_{p}(z)\bm{\Sigma}^{\frac{1}{2}}&\bm{\Sigma}^{\frac{1}{2}}\bar{\bm{Q}}_{p}(z)\bm{\Sigma}^{\frac{1}{2}}(\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}&\bm{0}\\ (\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}\bm{\Sigma}^{\frac{1}{2}}\bar{\bm{Q}}_{p}(z)\bm{\Sigma}^{\frac{1}{2}}&\frac{1}{z}\left((\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}\bm{\Sigma}^{\frac{1}{2}}\bar{\bm{Q}}_{p}(z)\bm{\Sigma}^{\frac{1}{2}}(\bm{R}^{{}^{\prime\prime}})^{\frac{1}{2}}-\bm{I}_{p}\right)&\bm{0}\\ \bm{0}&\bm{0}&\tilde{m}(z)\bm{I}_{n}\end{array}\right],$ (25)
where $\tilde{m}(z)=-z^{-1}(1+\phi^{1/2}e(z))^{-1}$. The anisotropic local law aims to control the difference $\bm{G}(z)-\bar{\bm{G}}(z)$ for $z\in\mathcal{D}$. A crucial observation is that defining the parameter $z$-dependent covariance $\bm{\Sigma}(z)=z\bm{\Sigma}(z\bm{I}_{p}-\bm{R}^{\prime})^{-1}$ with the corresponding measure $\pi^{z}=\sum_{i=1}^{p}\delta_{\lambda_{i}(\bm{\Sigma}(z))}/p$, where $\delta_{z}$ is the Dirac point measure at $z$, we have the following $z$-dependent deformed Marčenko-Pastur law,
$\frac{1}{\tilde{m}(z)}=-z+\int\frac{\phi^{\frac{1}{2}}t}{1+\phi^{-\frac{1}{2}}\tilde{m}(z)t}\mathrm{d}\pi^{z}(t).$ (26)
This result mirrors the form of the deformed Marčenko-Pastur law in Ding and Wang (2023), with $\bm{\Sigma}(z)$ as the covariance matrix. This demonstrates that the effect of $\bm{R}$ can be represented by turning $\bm{\Sigma}$ into the $z$-dependent covariance $\bm{\Sigma}(z)$. This insight simplifies our proof and presentation. For the denominator in (26), we impose the following technical assumption.
@@ 3.2 Universality
When $\alpha=1,$ we require that there exists $\tau>0$ such that

### v2

@@ 2 Universal bootstrap
The empirical distribution of $T^{\text{ub},b}$ for $b=1,\dots,B$ serves as an approximation for the distribution of $T^{\text{ub}}$. Specifically, we use the empirical upper-$\alpha$ quantile, $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}_{0}}(\alpha)$, of $T^{\text{ub},b}$, $b=1,\dots,B$, as the bootstrap approximated upper-$\alpha$ quantile of $T$. In particular, when testing $H_{0}:\bm{\Sigma}=\bm{\Sigma}_{0}$, we will reject $H_{0}$ if $T>\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0},\bm{\Sigma}_{0}}(\alpha)$, which does not contain unknown quantities. The complete universal bootstrap algorithm is presented in Algorithm 1.
input : Covariance matrix $\bm{\Sigma}$, $\bm{\Sigma}_{0}$, sample size $n$, test level $\alpha$, bootstrap number $B$
output : Bootstrapped $\alpha$ upper-quantile $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}_{0}}(\alpha)$
for $b\leftarrow 1$ to $B$ do
       $\bm{Z}^{b}\leftarrow$ $n\times p$ matrix with i.i.d. standard normal entries ;
       $\bm{Y}^{b}\leftarrow\bm{Z}^{b}\bm{\Sigma}^{1/2}$ ;
       $\bm{\hat{\Sigma}}^{\text{ub},b}\leftarrow(\bm{Y}^{b})^{T}(\bm{Y}^{b})/n$ ;
       $T^{\text{ub},b}\leftarrow\left\|\bm{\hat{\Sigma}}^{\text{ub},b}-\bm{\Sigma}_{0}\right\|_{\text{op}}$ ;
end for
$\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}_{0}}(\alpha)\leftarrow$ empirical upper-$\alpha$ quantile of $T^{\text{ub},b}$, $b=1,\dots,B$.
Algorithm 1 Universal bootstrap
We now extend the universal bootstrap procedure for $T$ to general spectral statistics for the covariance test. Similar to $T$ and $T^{\text{Roy}}$, such general spectral statistics can be constructed using the extreme singular values of $\bm{\Sigma}_{2}^{-1/2}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1})\bm{\Sigma}_{2}^{-1/2}$ for symmetric matrix $\bm{\Sigma}_{1}$ and positive-definite matrix $\bm{\Sigma}_{2}$. This inspires us to consider the statistic $T^{\text{ExS}}$ as a more general function of extreme singular values
$T^{\text{ExS}}=f\bigg{(}\bigg{\{}\left[\sigma_{k}\left(\bm{\Sigma}_{2,1}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,1})\bm{\Sigma}_{2,1}^{-\frac{1}{2}}\right)\right]_{k=1}^{k_{1}},\cdots,\left[\sigma_{k}\left(\bm{\Sigma}_{2,M}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}-\bm{\Sigma}_{1,M})\bm{\Sigma}_{2,M}^{-\frac{1}{2}}\right)\right]_{k=1}^{k_{M}}\bigg{\}}\bigg{)},$ (4)
where $f$ is a measurable function, $(\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m})$ are matrices, and $k_{m}$ are integers for $m=1,\cdots,M$. The statistic $T^{\text{ExS}}$ forms a general class of spectral statistics based on extreme singular values. For example, when $M=1$, $k_{1}=1$, $\bm{\Sigma}_{2,1}=\bm{I}_{p}$ and $\bm{\Sigma}_{1,1}=\bm{\Sigma}_{0}$, the statistic reduces to $T$. When $M=1$, $k_{1}=1$, $\bm{\Sigma}_{2,1}=\bm{\Sigma}_{0}$ and $\bm{\Sigma}_{1,1}=\bm{I}_{p}$, it reduces to $T^{\text{Roy}}$. This class also accommodates combinations of $T$ and $T^{\text{Roy}}$, regardless of their complex dependence. As discussed in Section 4, $T^{\text{Roy}}$ is more effective in worst-case scenarios, while $T$ performs better on average. To enhance testing power, we propose the following combined statistic
@@
This statistic belongs to the general class of $T^{\text{ExS}}$. Furthermore, while $T$ and $T^{\text{Roy}}$ rely only on the largest singular value (equivalent to the operator norm), $T^{\text{ExS}}$ incorporates the largest $k_{m}$ singular values, potentially improving test power Ke (2016).
The universal bootstrap extends naturally to the class of spectral statistics $T^{\text{ExS}}$. The universal bootstrapped version, $T^{\text{ExS,ub}}$, is obtained by replacing $\bm{\hat{\Sigma}}$ with $\bm{\hat{\Sigma}}^{\text{ub}}$ in (4). The empirical universal bootstrap threshold $\hat{q}^{\text{ExS,ub,B}}_{\bm{\Sigma},\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M}}(\alpha)$, is defined as the upper $\alpha$-th quantile of $T^{\text{ExS,ub,1}},\cdots,T^{\text{ExS,ub,B}}$ of the i.i.d. samples $T^{\text{ExS,ub}}$. The algorithm for this generalized universal bootstrap is provided in Algorithm 2. In Section 3, we establish the consistency of the universal bootstrap for generalized spectral statistics, demonstrating its applicability for constructing statistics using extreme eigenvalues in various ways.
input : Statistic function $f$, covariance matrix $\bm{\Sigma}$, $\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M}$ with number $k_{1},\cdots,k_{M}$, sample size $n$, test level $\alpha$, bootstrap number $B$
output : Bootstrapped $\alpha$ upper-quantile $\hat{q}^{\text{ExS,ub,B}}_{\bm{\Sigma},\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M}}(\alpha)$
for $b\leftarrow 1$ to $B$ do
       $\bm{Z}^{b}\leftarrow$ $n\times p$ matrix with i.i.d. standard normal entries ;
       $\bm{Y}^{b}\leftarrow\bm{Z}^{b}\bm{\Sigma}^{1/2}$ ;
       $\bm{\hat{\Sigma}}^{\text{ub},b}\leftarrow(\bm{Y}^{b})^{T}(\bm{Y}^{b})/n$ ;
@@
       $T^{\text{ub},b}\leftarrow f\bigg{(}\bigg{\{}\left[\sigma_{k}\left(\bm{\Sigma}_{2,1}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}^{\text{ub},b}-\bm{\Sigma}_{1,1})\bm{\Sigma}_{2,1}^{-\frac{1}{2}}\right)\right]_{k=1}^{k_{1}},\cdots,\left[\sigma_{k}\left(\bm{\Sigma}_{2,M}^{-\frac{1}{2}}(\bm{\hat{\Sigma}}^{\text{ub},b}-\bm{\Sigma}_{1,M})\bm{\Sigma}_{2,M}^{-\frac{1}{2}}\right)\right]_{k=1}^{k_{M}}\bigg{\}}\bigg{)}$;
end for
$\hat{q}^{\text{ExS,ub,B}}_{\bm{\Sigma},\left\{\bm{\Sigma}_{1,m},\bm{\Sigma}_{2,m}\right\}_{m=1}^{M}}(\alpha)\leftarrow$ empirical upper-$\alpha$ quantile of $T^{\text{ExS,ub,b}}$, $b=1,\dots,B$.
Algorithm 2 Generalized universal bootstrap
3 Theoretical guarantees
3.1 Universality properties
Our bootstrap consistency results rely on the universality from the random matrix theory. In this section, we review the background of random matrix theory and establish the universality property. To proceed, we first introduce some assumptions. We assume there exists some constants $C_{1},C_{2}$ and $\gamma\geq 1$ such that
$C_{1}n^{\gamma}\leq p\leq C_{2}n^{\gamma}.$ (6)
When $\gamma=1$, this is the typical proportional regime. When $\gamma>1$, the dimension-to-sample size ratio $p/n$ can diverge. The primary matrix of interest is $\bm{\hat{\Sigma}}-\bm{\Sigma}_{0}$. This inspires us to consider the matrix of the form $\bm{M}_{n}=\bm{\hat{\Sigma}}+\bm{R}$. Here, $\bm{R}$ is a symmetric matrix that may have positive, negative, or zero eigenvalues. We impose some assumptions below.
@@ 3.1 Universality properties
Suppose that $\bm{X}_{i}=\bm{\Sigma}^{1/2}\bm{Z}_{i}$ and $\bm{Z}_{i}=(Z_{i1},\cdots,Z_{ip})^{T}$, and $Z_{ij}$, $i=1,\cdots,n$, $j=1,\cdots,p$ are independent with $\mathbb{E}[Z_{ij}]=0$, $\mathbb{E}[Z_{ij}^{2}]=1$. There exists a positive sequence $C_{k}$ such that $\mathbb{E}[|Z_{ij}|^{k}]\leq C_{k}$ for $i=1,\cdots,n;j=1,\cdots,p$ and $k\in\mathbb{N}$.
@@
Assumptions 1 and 2 are standard assumptions in the random matrix literature (Alex et al. (2014)). Notably, Assumption 1 relies only on independence, without assuming identical distribution as in Qiu et al. (2023). While all moments exist under Assumption 1, this condition could be relaxed (Ding & Yang (2018)). We do not pursue this here. We also permit $\bm{\Sigma}$ to be singular, which is less restrictive than the invertibility assumed in Ding & Wang (2023); Ding et al. (2024). Lastly, Assumption 3, imposed for technical requirements, necessitates that $\bm{\Sigma}$ and $\bm{R}$ share the same eigenvectors. The same assumption is required in the signal-plus-noise model as in Zhou et al. (2023, 2024).
We introduce a parameter $z$-dependent covariance $\bm{\Sigma}(z)=z\bm{\Sigma}(z\bm{I}_{p}-\phi^{-1/2}\bm{R})^{-1}$ and $\tilde{m}(z)$ as the fixed point of the following $z$-dependent deformed Marčenko-Pastur law,
$\frac{1}{\tilde{m}(z)}=-z+\frac{1}{p}\sum_{i=1}^{p}\frac{\phi^{\frac{1}{2}}\lambda_{i}(\bm{\Sigma}(z))}{1+\phi^{-\frac{1}{2}}\tilde{m}(z)\lambda_{i}(\bm{\Sigma}(z))}.$ (7)
This result mirrors the deformed Marčenko-Pastur law in Ding & Wang (2023), with $\bm{\Sigma}(z)$ replacing the covariance matrix. The uniqueness and existence of the solution to (7) are provided in the supplementary material. Define $E_{+}$ and $E_{-}$ to be the deterministic limit of $\lambda_{1}(\phi^{-1/2}\bm{M}_{n})$ and $\lambda_{p}(\phi^{-1/2}\bm{M}_{n})$, respectively, whose formal definition is provided in the supplementary material. For the denominator in (7), we impose the following technical assumption.
@@
When $\gamma=1$ in (6), we require that there exists $\tau>0$ such that